[Brahms-l] Shift report 20040426 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Mon Apr 26 2004 - 23:43:03 EDT
------------------------------------------------------------
Date: Mon Apr 26 23:42:59 2004, Shift: 16:00-24:00
Supervisor: EBJ, with: 
Log book#15 pages 113 - 113

Shift Summary:
------------------------------------------------------------
1.  We had beam for most of the shift, but RHIC delivered stable beam at 7pm.

2. I completed 1/2 B 4 deg data set.  
   I have taken 130k of the 1/1 B 3 deg data set at the             
    moment.

3. I noticed that the gas flow rates were off, and after 
     taking to Ramiro.  I corrected them.  
   I also noticed that DVM3 was giving a funny distribution.  
     Ramiro tried to fix the problem from home, but I got 
     the impression that he wasn't sure that he fixed it.  
     Reguardless, the DVM3 supermon plot looks great now.    
     See the Elog.   
   Supermon has given me problems all evening.  It kept on 
      crashing at first, and now it is not printing all the 
      plots.  
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 26 23:43:16 2004

This archive was generated by hypermail 2.1.8 : Mon Apr 26 2004 - 23:43:38 EDT